# CU-JADE: A Method for Traversing Extinction Jumps along the Line-of-Sight

### v1

@@ Abstract
Although interstellar dust extinction serves as a powerful distance estimator, the solar system’s location within the Galactic plane complicates distance determinations, especially for molecular clouds (MCs) at varying distances along the line of sight (LoS). The presence of complex extinction patterns along LoS introduces degeneracies, resulting in less accurate distance measurements to overlapping MCs in crowded regions of the Galactic plane. In this study, we develop the CUSUM-based Jump-point Analysis for Distance Estimation (CU-JADE), a novel method designed to help mitigate these observational challenges. The key strengths of CU-JADE include: (1) sensitivity to detect abrupt jumps in $Distance$-$A_{\lambda}$ ($D$-$A$) datasets, (2) minimal systematic errors as demonstrated on both mock and observed data, and (3) the ability to combine CUSUM analysis with multi-wavelength data to improve the completeness of distance measurements for nearby gas structures, even for extinction values as low as $\Delta A_{V}\gtrsim 0.15$ mag. By combining CO survey data with a large sample of stars characterized by high-precision parallaxes and extinctions, we uncovered the multi-layered molecular gas distribution in the high-latitude Cepheus region. We also determined accurate distances to MCs beyond the Cygnus Rift by analyzing the intricate structure of gas and extinction within the Galactic plane. Additionally, we constructed a full-sky 3D extinction map extending to 4 kpc, which provides critical insights into dense interstellar medium components dominated by molecular hydrogen. These results advance our understanding of the spatial distribution and physical properties of MCs across the Milky Way.
@@ 1 Introduction
The selective absorption of starlight by the interstellar dust and gas leads to the dimming and reddening of starlight (Frerking et al., 1982; Schlegel et al., 1998). Utilizing the large star samples with high precision parallaxes and extinctions, we can investigate the characteristics of dust absorption in the infrared and visible spectrum. MCs traced by CO are typically accompanied by high extinction (mixed with a significant quantity of dust grains), yet the presence of dust extinction is not always correlated with CO-bright MCs (Mathis et al., 1983; Dame et al., 2001). The dust distribution exhibits a wide dynamic range from very dense cores to nearly transparent regions (Draine, 2003; Pineda et al., 2008; Leroy et al., 2011). And in low column density regions, significant dust emission can be observed in the mid-infrared, as well as the reddening effects in the optical and ultraviolet (Draine, 2011; Planck Collaboration et al., 2020; Shull et al., 2021). Although CO emission is very weak or even absent in low column density regions, $\rm H_{2}$ may exist based on observations and models (Grenier et al., 2005; Wolfire et al., 2010; Kalberla et al., 2020; Skalidis et al., 2024; Luo et al., 2024). We thus may use the extinction features to trace CO-faint or even CO-dark MCs.
The detailed three dimensional (3D) distribution and kinematic properties resolved along the Line of Sight (LoS) are essential for the study of MCs. For example, the superposition of crowded molecular gas components along the LoS directly impacts the mass-size relation of MCs (e.g., Ballesteros-Paredes et al., 2019; Lada & Dame, 2020), leading to misinterpretation of the indices derived from various statistical samples. And by examining a sample of stars within a specific distance range, some studies show that the polarization patterns of the ISM do not closely match the orientation of the magnetic field, which may indicate that the observations are affected by the superposition of multiple gas layers (Rezaei Kh. et al., 2020; Pelgrims et al., 2024; Panopoulou et al., 2024).
@@
Meanwhile, extinction jump methods, which employ single-extinction-jump models, are often applicable to high latitude regions (Schlafly et al., 2014; Yan et al., 2019b; Zucker et al., 2019; Zhao et al., 2020; Guo et al., 2022). And based on the identification of the MCs’ silhouette, which highlights the target MCs with selection of on-cloud stars and field stars, some works investigate the distance measurements for clouds near Galactic plane (Yan et al., 2019a; Chen et al., 2020; Zhang et al., 2024). However, the identification of MCs depends on the algorithms and changes with it. And on the other hand, in the inner Galactic region, the crowded MCs along the LoS cause significant extinction, leading to difficulties in detecting jumps in some more distant MCs and/or MCs with weak emission.
@@
The paper is organized as follows. In Section 2, we introduce the data we used in this work. Then we describe how the CU-JADE method works based on the characteristics of the LoS stellar parameters in Section 3, followed by validation of the method using mock and observed data in Section 4. In Section 5, we further implement the method on multilayer gas distance measurements for two typical regions toward the Cepheus and the Cygnus. We also present a 3D dust map extends to $\sim 4$ kpc for whole sky. Finally, in Section 6, we summarize the main results and conclusions.
@@ 2.1.1 MWISP Data
[…]
@@
The 1.2 m CfA survey (Dame et al., 2001) incorporates data from 37 individual $\rm{}^{12}CO$ ($J$ = 1–0) line surveys covering most of the Galactic plane. With an angular resolution of $68^{\prime}.$6, the surveys provide exceptional data homogeneity. Notably, it offers velocity coverage of $\pm$ 47.1 $\rm km~{}s^{-1}$ and significantly improved sensitivity of 0.18 K in 0.65 $\rm km~{}s^{-1}$ channels for high-latitude northern sky (Dame & Thaddeus, 2022) compared to the MWISP survey. We utilized the CfA dataset to systematically investigate dust-gas correlations in the Cepheus region, particularly focusing on the relationship between CO structures and the corresponding extinction features in $100^{\circ}\leq l\leq 120^{\circ}$, and $0^{\circ}\leq b\leq 20^{\circ}$.
@@ 2.2.1 Gaia DR3
Gaia’s Data Release 3 (i.e., Gaia DR3; Gaia Collaboration et al., 2023; Andrae et al., 2023) used the General Stellar Parameterizer from Photometry (GSP-Phot) to fit the BP/RP spectrum, parallax, and apparent $G$ magnitude simultaneously with a Bayesian forward-modelling approach. Over 470 million sources with $G<19$ were released. The $G$ band ($\lambda$ centered at 673 nm) extinction $A_{G}$ was constrained to be positive with roughly an error $\sim 0.06$ mag. For the Cygnus region near the Galactic plane, we retrieved the same stellar sample from Gaia DR3 and performed a comparative analysis in line with previous results (Zhang et al., 2024). Additionally, the statistical mock data on parallaxes and extinction discussed in Section 4 are also derived from this sample.
@@ 2.2.3 ZGR23
To effectively compare different extinction estimation methodologies (see details in Section 5.1.2), we adopted the comprehensive stellar catalog (hereafter ZGR23) published by Zhang et al. (2023), which provides astrophysical parameters for 220 million stars. This catalog was constructed using a forward modeling framework, which was trained on a carefully curated subset of high-resolution stellar spectra from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST), featuring a spectral resolution of $R\sim 1800$. The construction process also integrated multi-wavelength datasets, including Gaia BP/RP low-resolution spectra ($R\sim$ 30–100), 2MASS near-infrared photometry ($J$, $H$, Ks bands), and unWISE mid-infrared photometry ($W$1, $W$2 bands) (Gaia Collaboration et al., 2023; Montegriffo et al., 2023; Schlafly et al., 2019; Skrutskie et al., 2006; Wright et al., 2010). The derived parameters encompass critical astrophysical quanlities, such as distance estimates ($1/\varpi$; parallax-based), interstellar extinctions ($A_{V}$), and fundamental stellar properties, $\Theta\equiv(\text{T}_{\text{\rm eff}},~{}\log g,~{}[\text{Fe}/\text{H}])$.
[…]
@@ 2.3 Masers
To verify the accuracy of our method, we need MCs with precise distance measurements. Distances derived from observations of masers using Very Long Baseline Interferometry (VLBI, Xu et al., 2006; Reid et al., 2014) are considered to be among the most reliable measurements available. Class II masers (e.g., Claussen & Lo, 1986) can effectively trace the dense molecular gas clumps in high-mass star-forming regions (HMSFRs). In particular, many works (e.g., Reid et al., 2019; VERA Collaboration et al., 2020; Hyland et al., 2023) have measured the LoS velocities, proper motions, and trigonometric parallaxes of water and methanol masers associated with over 200 HMSFRs. We collected 75 masers with accurate parallax measurements within 3 kpc for validation in Section 4.2.
@@ 2.4 Young Open Clusters
We utilize young open clusters (YOCs; Hunt & Reffert, 2023) as tracers of star-forming regions. We select clusters with more than 30 stars and with distance uncertainty $<5$%. The age of the star cluster is limited to $log~{}Age~{}(yr)\lesssim$ 7 in order to establish the connection between the star-forming region and the corresponding gas-rich environment. That is, these YOCs might not move far from their birthplace, representing the dense molecular gas therein.
Figure 1: The flowchart illustrates the process of identifying extinction jump points in a $D$-$A$ diagram using the CU-JADE method. The process is demonstrated through a simple example with two jump points. Initialization: Begin with a set of simulated stellar samples exhibiting two jump points on the $D$-$A$ diagram (subplot a). These jump points are indicated by blue lines. CUSUM Statistic Calculation: Compute the CUSUM statistic $S$ to locate the first jump point (Step 1). Subplot b shows the CUSUM curve, where black points represent binned stellar samples at intervals of 20 pc, and the red curve represents the calculated CUSUM values using Equation (3). The position with the maximum $S$ value, marked by a red star, identifies the first jump point. Validity Check: Determine if the first jump point is valid (Step 2). This involves calculating the $S_{\rm diff}$ value and comparing it to the $S_{\rm diff}$ background (see Section 3.2 for details). Subplots c and d illustrate this comparison using percentile, similar to a confidence level, with a parameter $C$ for adjustment. Splitting and Recalculation: If the first jump point is confirmed, split the $D$-$A$ data at this position (Step 3, see subplots b and e). Recalculate the separate datasets and repeat the process, as indicated by the dashed arrow. This begins the traversal of all extinction jump points (suggested by bulges on the CUSUM curve). Final Jump Point Identification: Since only two jump points were initially set, the next round of calculations locates the last jump point (Step 4, subplot f). And the process enables us to return all detected jump point positions (Step 5). Validation: Compare and check the detected jump point positions against the preset positions (Step 6, subplots g and a).
@@ 3.1 The CUSUM Chart-based method
The increment of LoS extinctions is statistically discernible as a multiple jumps in Gaussian-Mean problem (Andrae et al., 2018; Fearnhead & Fryzlewicz, 2024). The CUSUM chart is known for its sensitivity to small shifts within a series of samples, thereby enabling the detection of jumps in star extinction features caused by the ISM. For the first time, our study applies CU-JADE method to identify extinction jump points along the $Distance$-$A_{\lambda}$ ($D$-$A$) data sequence.
In this section, we adopt the techniques in Taylor (2000), and improve them to suit our scenario. Considering a sequence of $N$ star samples exhibiting two distinct extinction jump points in the $D$-$A$ space (Figure 1a), denoted as $A_{\lambda}$: $\{A_{1},A_{2},\cdots,A_{i},\cdots,A_{N}\}$, along with the corresponding distance of each star $D$: $\{D_{1},D_{2},\cdots,D_{i},\cdots,D_{N}\}$.
@@
Sort the $A_{\lambda}$ in ascending order based on distance. Compute the CUSUM statistic $S$ for the remaining ordered values.
@@
For any given series of $A_{\lambda}$ values, it returns a sequence of cumulative sum values in this manner: $\{S_{1},S_{2},\cdots,S_{i},\cdots,S_{N}\}$. In this case, for a group of $A_{\lambda}$ values in ascending order, the $Distance$-CUSUM $S$ ($D$-$S$) plot shows a trend where $S_{i}$ increases initially, and after reaching the peak, it descends until it drops to zero (see red curve in Figure 1b). The prominent peak (denoted by a red star in the CUSUM curve) is preliminarily identified as the first extinction jump point.
The turning point is caused by the change in the difference between the $A_{\lambda}$ values and their mean $\bar{A}$. In the absence of extinction jump points, or when the data sequence is randomly permuted, the CUSUM statistic $S$ would exhibit random fluctuations, characterized by the absence of statistically significant peaks or bulges (see colored curves in Figure 1c). This feature can help us determine whether the identified jump point is valid.
@@
The confidence level ($CL$) regarding the existence of a jump point in the $D$-$A$ sequence can be assessed using $S_{\rm diff}$, which characterizes the prominence of the identified spike or turning point in the CUSUM analysis. This value is calculated as follows:
@@
To enhance the statistical robustness of the $CL$ and mitigate the risk of false positive detections, we introduce a scaling factor $C$ to modulate the statistic $S_{\rm diff}$: $\{C\cdot S_{\rm diff}^{1},C\cdot S_{\rm diff}^{2},\cdots,C\cdot S_{\rm diff}^{N_{\rm shuf}}\}$. The optimal $C$ value ($C=1-3$ is usually enough, $C=5$ in the Figure 1c and d) can be determined through validation testing, using both observational catalogs and mock data (see Section 4). Finally, the presence of a jump point is evaluated using a predefined $CL$ threshold, typically set at 0.95.
@@
During the iteration, from panel (b) to (f) in Figure 1, we find that the CUSUM statistic $S^{0}_{\rm diff}$ of an ordered sequence drops significantly. We attribute the progressive decline in $S^{0}_{\rm diff}$ during Step 3 to the variation in the jump value $\Delta A_{\lambda}$, the number of stellar sample, and the caveats from early-detected jump points. The last factor indicates that the first detected jump points are not necessarily the most significant ones and may carry elevated false-positive risks. Focusing on the final set of detected jumps (our termination condition) might simplify our understanding for interpreting multi-jump scenarios. In the next subsection, we quantitatively analyze factors influencing the $CL$ of jump point detection. For extended discussion and complications, see the Appendix A.1.
Step 5–6. Return all jump points and check against input/external constraints.

### v2

# CU-JADE: A Method for Traversing Extinction Jumps along the Line of Sight

@@ Abstract
Although interstellar dust extinction serves as a powerful distance estimator, the solar system’s location within the Galactic plane complicates distance determinations, especially for molecular clouds (MCs) at varying distances along the line of sight (LoS). The presence of complex extinction patterns along the LoS introduces degeneracies, resulting in less accurate distance measurements to overlapping MCs in crowded regions of the Galactic plane. In this study, we develop the CUSUM-based Jump-point Analysis for Distance Estimation (CU-JADE), a novel method designed to help mitigate these observational challenges. The key strengths of CU-JADE include: (1) sensitivity to detect abrupt jumps in Distance-$A_{\lambda}$ ($D$-AA) datasets, (2) minimal systematic errors as demonstrated on both mock and observed data, and (3) the ability to combine CUSUM analysis with multiwavelength data to improve the completeness of distance measurements for nearby gas structures, even for extinction values as low as $\Delta A_{V}\gtrsim 0.15$ mag. By combining CO survey data with a large sample of stars characterized by high-precision parallaxes and extinctions, we uncovered the multilayered molecular gas distribution in the high-latitude Cepheus region. We also determined accurate distances to MCs beyond the Cygnus Rift by analyzing the intricate structure of gas and extinction within the Galactic plane. Additionally, we constructed a full-sky 3D extinction map extending to 4 kpc, which provides critical insights into dense interstellar medium components dominated by molecular hydrogen. These results advance our understanding of the spatial distribution and physical properties of MCs across the Milky Way.
@@ 1 Introduction
The selective absorption of starlight by the interstellar dust and gas leads to the dimming and reddening of starlight (Frerking et al., 1982; Schlegel et al., 1998). Utilizing the large star samples with high precision parallaxes and extinctions, we can investigate the characteristics of dust absorption in the infrared and visible spectrum. MCs traced by CO are typically accompanied by high extinction (mixed with a significant quantity of dust grains), yet the presence of dust extinction is not always correlated with CO-bright MCs (Mathis et al., 1983; Dame et al., 2001). The dust distribution exhibits a wide dynamic range from very dense cores to nearly transparent regions (Draine, 2003; Pineda et al., 2008; Leroy et al., 2011). In addition, in low column density regions, significant dust emission can be observed in the mid-infrared, as well as the reddening effects in the optical and ultraviolet (Draine, 2011; Planck Collaboration et al., 2020; Shull et al., 2021). Although CO emission is very weak or even absent in low column density regions, $\rm H_{2}$ may exist based on observations and models (Grenier et al., 2005; Wolfire et al., 2010; Kalberla et al., 2020; Luo et al., 2024; Skalidis et al., 2024). We thus may use the extinction features to trace CO-faint or even CO-dark MCs.
The detailed 3D distribution and kinematic properties resolved along the line of sight (LoS) are essential for the study of MCs. For example, the superposition of crowded molecular gas components along the LoS directly impacts the mass-size relation of MCs (e.g., Ballesteros-Paredes et al., 2019; Lada & Dame, 2020), leading to misinterpretation of the indices derived from various statistical samples. By examining a sample of stars within a specific distance range, some studies show that the polarization patterns of the ISM do not closely match the orientation of the magnetic field, which may indicate that the observations are affected by the superposition of multiple gas layers (Rezaei Kh. et al., 2020; Pelgrims et al., 2024; Panopoulou et al., 2024).
@@
Meanwhile, extinction-jump methods, which employ single-extinction-jump models, are often applicable to high-latitude regions (Schlafly et al., 2014; Yan et al., 2019b; Zucker et al., 2019; Zhao et al., 2020; Guo et al., 2022). Based on the identification of the MCs’ silhouette, which highlights the target MCs with selection of on-cloud stars and field stars, some works investigate the distance measurements for clouds near Galactic plane (Yan et al., 2019a; Chen et al., 2020; Zhang et al., 2024). However, the identification of MCs depends on the algorithms and changes with it. On the other hand, in the inner Galactic region, the crowded MCs along the LoS cause significant extinction, leading to difficulties in detecting jumps in some more distant MCs and/or MCs with weak emission.
@@
This paper is organized as follows. In Section 2, we introduce the data we used in this work. Then, we describe how the CU-JADE method works based on the characteristics of the LoS stellar parameters in Section 3, followed by validation of the method using mock and observed data in Section 4. In Section 5, we further implement the method on multilayer gas distance measurements for two typical regions toward the Cepheus and the Cygnus. We also present a 3D dust map that extends to $\sim 4$ kpc for whole sky. Finally, in Section 6, we summarize the main results and conclusions.
@@ 2.1.1 MWISP Data
The CO data in the Galactic plane originates from the Milky Way Imaging Scroll Painting (MWISP project, for more details, refer to Su et al., 2019). In brief, the spatial resolution of the observational data is $\sim 50^{\prime\prime}$, and the spectral resolution is $\sim 0.2$ $\rm km~s^{-1}$. The processed 3D data cubes, referred to as position-position-velocity space, have a grid spacing of $30^{\prime\prime}$. The typical root mean square (rms) noise level is $\sim 0.5$ K for $\rm{}^{12}CO$ and $\sim 0.3$ K for $\rm{}^{13}CO$ and $\rm C^{18}O$. In this work, we focus on MCs located in a specific region of the Galactic plane toward the Cygnus region (i.e., $72^{\circ}\leq l\leq 87^{\circ}$, $|b|\leq 5^{\circ}.1$; Zhang et al., 2024).
@@
The 1.2 m CfA survey (Dame et al., 2001) incorporates data from 37 individual $\rm{}^{12}CO$ ($J$ = 1–0) line surveys covering most of the Galactic plane. With an angular resolution of $8^{\prime}.6$, the surveys provide exceptional data homogeneity. Notably, it offers velocity coverage of $\pm$ 47.1 $\rm km~s^{-1}$ and significantly improved sensitivity of 0.18 K in 0.65 $\rm km~s^{-1}$ channels for high-latitude northern sky (Dame & Thaddeus, 2022) compared to the MWISP survey. We utilized the CfA dataset to systematically investigate dust-gas correlations in the Cepheus region, particularly focusing on the relationship between CO structures and the corresponding extinction features in $100^{\circ}\leq l\leq 120^{\circ}$, and $0^{\circ}\leq b\leq 20^{\circ}$.
@@ 2.2.1 Gaia DR3
Gaia’s Data Release 3 (i.e., Gaia DR3; Gaia Collaboration et al., 2023; Andrae et al., 2023) used the General Stellar Parameterizer from Photometry (GSP-Phot) to fit the BP/RP spectrum, parallax, and apparent $G$ magnitude simultaneously with a Bayesian forward-modelling approach. Over 470 million sources with $G<19$ were released. The $G$-band ($\lambda$ centered at 673 nm) extinction $A_{G}$ was constrained to be positive with roughly an error $\sim 0.06$ mag. For the Cygnus region near the Galactic plane, we retrieved the same stellar sample from Gaia DR3 and performed a comparative analysis in line with previous results (Zhang et al., 2024). Additionally, the statistical mock data on parallaxes and extinction discussed in Section 4 are also derived from this sample.
@@ 2.2.3 ZGR23
To effectively compare different extinction estimation methodologies (see details in Section 5.1.2), we adopted the comprehensive stellar catalog (hereafter ZGR23) published by Zhang et al. (2023), which provides astrophysical parameters for 220 million stars. This catalog was constructed using a forward-modeling framework, which was trained on a carefully curated subset of high-resolution stellar spectra from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope, featuring a spectral resolution of $R\sim 1800$. The construction process also integrated multiwavelength datasets, including Gaia BP/RP low-resolution spectra ($R\sim$ 30–100), 2MASS near-infrared photometry ($J$, $H$, $K$s bands), and unWISE mid-infrared photometry (WW1, WW2 bands) (Gaia Collaboration et al., 2023; Montegriffo et al., 2023; Schlafly et al., 2019; Skrutskie et al., 2006; Wright et al., 2010). The derived parameters encompass critical astrophysical quantities, such as distance estimates ($1/\varpi$; parallax-based), interstellar extinctions ($A_{V}$), and fundamental stellar properties, $\Theta\equiv(\text{T}_{\text{\rm eff}},~\log g,~[\text{Fe}/\text{H}])$.
To generate the extinction map, we implemented additional section criteria to ensure data reliability and accuracy: 1. Quality Control: Sources were included only if their $\text{quality\_flags}<$ 8; 2. Distance Constraint: A distance cutoff of $1/\varpi<$ 5 kpc was applied; 3. Precision Threshold: A relative parallax precision threshold of $\sigma_{\varpi}/\varpi<$ 0.2 was imposed to minimize uncertainties in distance measurements. These criteria resulted in a refined sample of 123,284,327 stars, enabling accurate 3D extinction mapping and systematic comparison of different methods.
@@ 2.3 Masers
To verify the accuracy of our method, we need MCs with precise distance measurements. Distances derived from observations of masers using Very Long Baseline Interferometry (Xu et al., 2006; Reid et al., 2014) are considered to be among the most reliable measurements available. Class II masers (e.g., Claussen & Lo, 1986) can effectively trace the dense molecular gas clumps in high-mass star-forming regions (HMSFRs). In particular, many works (e.g., Reid et al., 2019; VERA Collaboration et al., 2020; Hyland et al., 2023) have measured the LoS velocities, proper motions, and trigonometric parallaxes of water and methanol masers associated with over 200 HMSFRs. We collected 75 masers with accurate parallax measurements within 3 kpc for validation in Section 4.2.
@@ 2.4 Young Open Clusters
We utilize young open clusters (YOCs; Hunt & Reffert, 2023) as tracers of star-forming regions. We select clusters with more than 30 stars and with distance uncertainty <5<5%. The age of the star cluster is limited to $log~Age~(yr)\lesssim$ 7 in order to establish the connection between the star-forming region and the corresponding gas-rich environment. That is, these YOCs might not move far from their birthplace, representing the dense molecular gas therein.
Figure 1: This flowchart illustrates the process of identifying extinction-jump points in a $D$-AA diagram using the CU-JADE method. The process is demonstrated through a simple example with two jump points. Initialization: Begin with a set of simulated stellar samples exhibiting two jump points on the $D$-AA diagram (subplot (a)). These jump points are indicated by blue lines. CUSUM Statistic Calculation: Compute the CUSUM statistic $S$ to locate the first jump point (Step 1). Subplot b shows the CUSUM curve, where black points represent binned stellar samples at intervals of 20 pc, and the red curve represents the calculated CUSUM values using Equation (3). The position with the maximum $S$ value, marked by a red star, identifies the first jump point. Validity check: Determine if the first jump point is valid (Step 2). This involves calculating the $S_{\rm diff}$ value and comparing it to the $S_{\rm diff}$ background (see Section 3.2 for details). Subplots (c) and (d) illustrate this comparison using percentile, similar to a confidence level, with a parameter $C$ for adjustment. Splitting and recalculation: If the first jump point is confirmed, split the $D$-AA data at this position (Step 3, see subplots (b) and (e)). Recalculate the separate datasets and repeat the process, as indicated by the dashed arrow. This begins the traversal of all extinction-jump points (suggested by bulges on the CUSUM curve). Final jump point identification: Since only two jump points were initially set, the next round of calculations locates the last jump point (Step 4, subplot (f)). And the process enables us to return all detected jump point positions (Step 5). Validation: Compare and check the detected jump point positions against the preset positions (Step 6, subplots (g) and (a)).
@@ 3.1 The CUSUM Chart-based method
The increment of LoS extinctions is statistically discernible as a multiple jumps in Gaussian-Mean problem (Andrae et al., 2018; Fearnhead & Fryzlewicz, 2024). The CUSUM chart is known for its sensitivity to small shifts within a series of samples, thereby enabling the detection of jumps in star extinction features caused by the ISM. For the first time, our study applies CU-JADE method to identify extinction-jump points along the Distance-$A_{\lambda}$ ($D$-AA) data sequence.
In this section, we adopt the techniques in Taylor (2000), and improve them to suit our scenario. Considering a sequence of $N$ star samples exhibiting two distinct extinction-jump points in the $D$-AA space (Figure 1a), denoted as $A_{\lambda}$: $\{A_{1},A_{2},\cdots,A_{i},\cdots,A_{N}\}$, along with the corresponding distance of each star $D$: $\{D_{1},D_{2},\cdots,D_{i},\cdots,D_{N}\}$.
@@
Sort the $A_{\lambda}$ in ascending order based on distance. Compute the CUSUM statistic $S$ for the remaining ordered values,
@@
For any given series of $A_{\lambda}$ values, it returns a sequence of cumulative sum values in this manner: $\{S_{1},S_{2},\cdots,S_{i},\cdots,S_{N}\}$. In this case, for a group of $A_{\lambda}$ values in ascending order, the Distance-CUSUM $S$ ($D$-$S$) plot shows a trend where $S_{i}$ increases initially, and after reaching the peak it descends until it drops to zero (see red curve in Figure 1b). The prominent peak (denoted by a red star in the CUSUM curve) is preliminarily identified as the first extinction-jump point.
The turning point is caused by the change in the difference between the $A_{\lambda}$ values and their mean $\bar{A}$. In the absence of extinction-jump points, or when the data sequence is randomly permuted, the CUSUM statistic $S$ would exhibit random fluctuations, characterized by the absence of statistically significant peaks or bulges (see colored curves in Figure 1c). This feature can help us determine whether the identified jump point is valid.
@@
The confidence level ($CL$) regarding the existence of a jump point in the $D$-AA sequence can be assessed using $S_{\rm diff}$, which characterizes the prominence of the identified spike or turning point in the CUSUM analysis. This value is calculated as follows:
@@
To enhance the statistical robustness of the $CL$ and mitigate the risk of false positive detections, we introduce a scaling factor $C$ to modulate the statistic $S_{\rm diff}$: $\{C\cdot S_{\rm diff}^{1},C\cdot S_{\rm diff}^{2},\cdots,C\cdot S_{\rm diff}^{N_{\rm shuf}}\}$. The optimal $C$ value (C=13C=1-3 is usually enough, $C=5$ in the Figure 1c and d) can be determined through validation testing, using both observational catalogs and mock data (see Section 4). Finally, the presence of a jump point is evaluated using a predefined $CL$ threshold, typically set at 0.95.
@@
During the iteration, from panel (b) to (f) in Figure 1, we find that the CUSUM statistic $S^{0}_{\rm diff}$ of an ordered sequence drops significantly. We attribute the progressive decline in $S^{0}_{\rm diff}$ during Step 3 to the variation in the jump value $\Delta A_{\lambda}$, the number of stellar sample, and the caveats from early-detected jump points. The last factor indicates that the first detected jump points are not necessarily the most significant ones and may carry elevated false-positive risks. Focusing on the final set of detected jumps (our termination condition) might simplify our understanding for interpreting multijump scenarios. In the next subsection, we quantitatively analyze factors influencing the $CL$ of jump point detection. For extended discussion and complications, see the Appendix A.1.
Steps 5–6. Return all jump points and check against input/external constraints.